# Conditioning diffusion models by explicit forward-backward bridging

### v1

@@ 3 Separable dynamics
with block diagonal $\Sigma_{t}$ and $A_{t}$. In this section, we describe how our approach can be further improved or adapted for these models.
@@ 3.1 Improving the efficiency of the CSMC
We note that this can be done approximately in the context of non-separable nonlinear SDEs (Bladt et al., 2016), for example, arising from Schrödinger bridges. However, since we specifically focus on exact conditioning, we elect not to treat this case here.
@@ 3.2 A pseudo-marginal implementation
When the forward noising process is given as in (7), we can implement a pseudo-marginal counterpart of Section 2, known as the particle Markov chain Monte Carlo (PMCMC) method, or more precisely the particle marginal Metropolis–Hastings (PMMH) algorithm (Andrieu et al., 2010). This approach relies on the fact that particle filtering returns an unbiased estimate of the marginal likelihood of the path $Y_{1:K}$ under $\mathbb{B}$, given as $\prod_{k=1}^{K}\frac{1}{N}\sum_{n=1}^{N}w_{k}^{n}$, outputted by Algorithm 1 (with $F=0$), which can be used in a Metropolis–Hastings step to sample from the posterior distribution of the path $X_{0:K}$. In the context of our problem, this corresponds to the following procedure. Given $X^{j}_{0}$ distributed according to $\pi(\cdot\mid y)$ and an unbiased estimator of the marginal likelihood $Z^{j}$ of $\mathbb{B}(y_{K:0})=\int\mathbb{B}(y_{K:0},\mathop{}\!\mathrm{d}u_{0:K})$: {mdframed}[style=enumFrame]
@@
In step 2, we take an $\mathbb{F}(\cdot\mid y_{0})$-invariant kernel. In this work, we choose the preconditioned Crank–Nicolson (PCN, Cotter et al., 2013) proposal, which is well adapted to Gaussian priors such as (7). Noting that $\mathbb{F}(y_{1:K}\mid y_{0}=y)$ is a linear transformation of Gaussian variables $\eta$, and writing $\eta^{j}$ for those corresponding to $Y^{j}_{1:K}$, this consists in sampling $q$ as follows: given a step size $\delta>0$, (i) sample $\eta^{\prime}\sim\mathcal{N}(0,I)$, (ii) set $\eta^{*}=\frac{2}{2+\delta}\,\eta^{j}+\sqrt{1-\frac{4}{4+4\delta+\delta^{2}}}\,\eta^{\prime}$, and (iii) form $Y^{*}$ from $\eta^{*}$. Importantly, this can be done with zero memory cost if one has access to the random number generator that was used to sample $\eta^{j}$ originally, in which case we can simply re-simulate $\eta^{j}$ alongside $\eta^{\prime}$ at zero memory cost and negligible (compared to evaluating the SDE drift) computational cost.
@@ 4 Experiments
In this section, we conduct quantitative and qualitative experiments to validate our proposed method (i.e., Gibbs-CSMC in Sections 2.1 and 3.1 and PMCMC in Section 3.2) on both synthetic and real data. We compare to the methods most related to ours: the standard particle filter (PF, Trippe et al., 2023; Dou and Song, 2024), twisted particle filter (TPF, Wu et al., 2023), and conditional score matching (CSGM, Song et al., 2021). Throughout our experiments, when a particle filter is used, we implement it using stratified resampling (see, e.g., Chopin and Papaspiliopoulos, 2020, Chap. 9), and when we use a conditional particle filter, we implement it using the conditional “killing” resampling of Karppinen et al. (2023), rather than the multinomial resampling as described in Algorithm 1. Unless otherwise mentioned, we repeat every experiment 100 times independently, and we report the statistics (i.e., mean, standard deviation, and/or quantile) of their results. For ease of reproducibility, we publish our implementations at https://github.com/zgbkdlm/fbs.
4.1 High-dimensional synthetic conditional sampling
In our first experiment, we assess the convergence of all the conditioning methods considered to the ‘true’ conditional distribution. To do so, we set the target $\pi(x\mid y)$ as the posterior distribution of a Gaussian process (GP) regression model for which the ground truth is available. We choose a standard Ornstein–Uhlenbeck forward noising process, for which we can compute the reverse SDE (2) exactly, removing the discretisation bias to isolate the error coming from the conditioning only. For each method, we generate 10,000 posterior samples, and then compute how these samples approximate the true posterior distribution in terms of Kullback–Leibler (KL) divergence, Wasserstein–Bures (Bures) distance, and mean absolute errors of the marginal means and variances. The detailed experimental setting is given in Appendix B.2.
Table 1: Errors statistics for the the conditional sampling problem of Section 4.1. The left and right panels show the errors when using 10 and 100 particles, respectively. The columns “Mean” and “Variance” report the errors in the marginal mean and variance of the posterior distribution. The number after PMCMC is the PCN parameter $\delta$ applied. Note that the CSGM approach can be computed exactly, we therefore use it as the reference and do not compete against it here.
@@ 4.1 High-dimensional synthetic conditional sampling
As demonstrated by Table 4.1, Gibbs-CSMC significantly outperforms all other methods for all metrics. Importantly, the Gibbs-CSMC method is substantially better than PF, showing that our MCMC strategy effectively corrects the bias of the PF approach of Trippe et al. (2023). Our PMCMC approach also improves upon PF but is sensitive to the calibration of the PCN parameter $\delta$. Specifically, when $\delta=0.005$, the KL of PMCMC is the worst, and decreasing it to $\delta=0.001$ improves the errors. On the other hand, the marginal mean and variance errors of PMCMC are relatively good, compared to KL, indicating that the method does not approximate well the off-diagonal part of the GP posterior covariance. The results of TPF appear to be worse than PF, and also have larger standard deviations. Finally, we also see from the table that increasing the number of particles improves the errors for all methods. More qualitative results are given in Appendix B.2.
Figure 1: Autocorrelations of PMCMC and Gibbs-CSMC in Section 4.1. For each of the 100 dimensions we compute the autocorrelation averaged over four independent chains. We then report the autocorrelation of the worst dimension. We see that Gibbs-CSMC is better than PMCMC, and that increasing the number of particles improves the autocorrelation.
Figure 2: Conditional sampling errors on a Gaussian Schrödinger bridge. PF (ideal) refers to using the exact posterior distribution (which is not available in general) to initialise PF, whereas PF (approximate) uses a standard Normal at initialisation. Gibbs sampling improves the sampling quality even compared to PF (ideal), highlighting the incorrectness of PF.
Figure 2 plots the autocorrelations of our Gibbs-CSMC and PMCMC methods. We see that the autocorrelation of Gibbs-CSMC is better than PMCMC, even when using a lower number of particles. We also observe that the autocorrelation of PMCMC can be improved by decreasing the PCN parameter $\delta$. This is mostly due to the variance of the log-likelihood estimation (see Section 3.2) of the PMCMC method (Andrieu and Roberts, 2009; Doucet et al., 2015), sometimes preventing the algorithm from moving, and which is negatively affected by $\delta$. Lastly, the figure also confirms that increasing the number of particles improves the autocorrelations overall.
@@ 4.2 Conditional sampling on non-separable noising processes
One key merit of our method is its generalisability: it works on non-separable noising processes as well as separable ones. Precisely, our method does not need $X_{t}$ and $Y_{t}$ of the noising process in (4) to be independent conditionally on $X_{0}$ and $Y_{0}$. On the other hand, the separability is a requirement for the standard PF approach of Dou and Song (2024); Trippe et al. (2023). Indeed, when the noising process is non-separable, their approach requires us to sample $X_{0}$ in order to generate a path $t\mapsto Y_{t}$. Ideally, $X_{0}\sim\pi(x\mid y)$ should come from the true conditional distribution, but in reality, we have to use approximate conditional samplers, which incur additional biases that are hard to eliminate.
To quantify the bias arising from the non-separability, we perform conditional sampling on a Gaussian Schrödinger bridge model (for which closed-form expressions are available, see, e.g., Bunne et al., 2023) using our Gibbs-CSMC approach and the PF of Trippe et al. (2023). We do not compare to our PMCMC approach given its non-applicability here (see Section 3.2). We continue working on the same GP regression problem in Section 4.1, but, to highlight further the impact of non-separability, rather than using a reference $\mathcal{N}(0,I)$, we use $\pi_{T}\sim\mathcal{N}(0,\Sigma)$, where $\Sigma$ is a random covariance matrix drawn from a Wishart distribution. The results are shown in Figure 2. We see that approximating the initial $\pi(x\mid y)$ introduces significant errors compared to the PF using exact initial sampling. Moreover, increasing the number of particles does not compensate for this initial bias highlighting the defect of the standard approach in non-separable models.
Figure 3: MNIST super-resolution (x4) on a (non-separable) Schrödinger bridge noising process. For each method, we show two samples (more samples are shown in Figure 7). We find that PF is significantly affected by its initialisation of $X_{0}$, while our Gibbs-CSMC method barely suffers from this problem.
We also qualitatively show the results for MNIST super-resolution on a pre-trained Schrödinger bridge. This time, we no longer have access to an exact $X_{0}$ sampler for the PF method, and we here use a uniform random $X_{0}$, and an $X_{0}$ using linear interpolation of $y_{0}$. The results are shown in Figure 3. We see from this figure that the initial specification of $X_{0}$ appearing in the forward noising process significantly impacts the results. When using a random $X_{0}$, where all pixels are uniform $\mathcal{U}(0,1)$, the recovered images deviate significantly from the truth. When using linear interpolation, which is a better initialisation, the recovered images look closer to the truth, but the quality is still poor. On the other hand, Gibbs-CSMC does not suffer from this initialisation problem and recovers the original digit as the MCMC chain converges. More results are given in Appendix B.3.
@@ 4.3 Inpainting and super-resolution
In this section, we apply our methods for inpainting and super-resolution on MNIST (resolution 28 by 28) and CelebA-HQ (resolution 64 by 64) images. For a complete description of the inpainting and super-resolution problems and additional experimental details, we refer the reader to Appendix B.4.
Given pre-trained (via the standard score matching method of Song et al., 2021) unconditional generative diffusions for MNIST and Celeba-HQ images, we then perform conditional sampling of missing (inpainting) or deblurred (super-resolution) of images given an observation thereof. To quantify the results, for each test image we recover 100 conditional image samples, and we compute the peak signal-to-noise ratio (PSNR), structural similarity index (SSIM), and learnt perceptual patch similarity (LPIPS, Zhang et al., 2018) metrics averaged over the 100 samples.
@@
PF $15.85$ $0.74$ $10.14$ $0.32$ $17.34$ $0.79$ $11.66$ $0.45$
Gibbs-CSMC $16.79$ $0.77$ $10.84$ $0.38$ $17.82$ $0.82$ $14.06$ $0.61$
PMCMC-0.005 $\mathbf{17.51}$ $\mathbf{0.80}$ $\mathbf{12.88}$ $\mathbf{0.54}$ $\mathbf{17.84}$ $\mathbf{0.81}$ $\mathbf{14.22}$ $\mathbf{0.63}$
TPF $13.45$ $0.48$ $9.67$ $0.19$ $13.89$ $0.51$ $10.38$ $0.25$
CSGM $15.17$ $0.71$ $9.72$ $0.28$leftarrow $\leftarrow$ $\leftarrow$ $\leftarrow$
@@
PF 22.88 0.85 0.05 23.29 0.79 0.11 23.98 0.87 0.04 24.57 0.82 0.09
Gibbs-CSMC 22.86 0.86 0.05 23.75 0.80 0.11 24.22 0.87 0.04 25.07 0.84 0.09
PMCMC-0.005 23.96 0.87 0.04 24.37 0.82 0.10 24.68 0.88 0.04 25.30 0.85 0.07
TPF 13.79 0.59 0.20 12.71 0.44 0.23 13.93 0.58 0.20 12.92 0.45 0.23
CSGM 22.86 0.85 0.05 23.48 0.79 0.11 $\leftarrow$ $\leftarrow$ $\leftarrow$ $\leftarrow$ $\leftarrow$ $\leftarrow$
@@
In Figure 4, we illustrate the methods for inpainting and super-resolution on a single task. We see that PF, TPF, and CSGM can generate unrealistic samples, and the super-resolution samples are sometimes contaminated by pixel distortions (e.g., the isolated white dots in the second panel, first row, third-fifth columns in the figure). This is particularly true for TPF despite it largely higher computational cost (due to taking gradient with respect to the score function in its twisting function, Wu et al., 2023) as it requires to further approximate the likelihood model $\pi(y\mid x)$ by a smoother version to be usable.
However, our Gibbs-CSMC and PMCMC methods tend to produce better quality samples, albeit more correlated ones, as those wrong samples are rejected by the MCMC kernel. For more examples and results, we refer the readers to Appendix B.4.

### v2

@@ 3 SEPARABLE DYNAMICS
with block diagonal $\Sigma_{t}$ and $A_{t}$. This is particularly favourable for image inpainting problems to leverage pre-trained diffusion models, where $x$ and $y$ stand for the missing and observed parts, repectively, of an image. In this section, we describe how our approach can be further improved or adapted for these models.
@@ 3.1 IMPROVING THE EFFICIENCY OF THE CSMC
We note that this can be done approximately in the context of non-separable nonlinear SDEs (Bladt et al.,, 2016), for example, arising from Schrödinger bridges. However, since we specifically focus on exact conditioning, we elect not to treat this case here.
@@ 3.2 A PSEUDO-MARGINAL IMPLEMENTATION
When the forward noising process is given as in (7), we can implement a pseudo-marginal counterpart of Section 2, known as the particle Markov chain Monte Carlo (PMCMC) method, or more precisely the particle marginal Metropolis–Hastings (PMMH) algorithm (Andrieu et al.,, 2010). This approach relies on the fact that particle filtering returns an unbiased estimate of the marginal likelihood of the path $Y_{1:K}$ under $\mathbb{B}$, given as $\prod_{k=1}^{K}\frac{1}{N}\sum_{n=1}^{N}w_{k}^{n}$, outputted by Algorithm 1 (with $F=0$), which can be used in a Metropolis–Hastings step to sample from the posterior distribution of the path $X_{0:K}$. In the context of our problem, this corresponds to the following procedure. Given $X^{j}_{0}$ distributed according to $\pi(\cdot\mid y)$ and an unbiased estimator of the marginal likelihood $Z^{j}$ of $\mathbb{B}(y_{K:0})=\int\mathbb{B}(y_{K:0},\mathop{}\!\mathrm{d}u_{0:K})$: {mdframed}[style=enumFrame]
@@
In step 2, we take an $\mathbb{F}(\cdot\mid y_{0})$-invariant kernel. In this work, we choose the preconditioned Crank–Nicolson (PCN, Cotter et al.,, 2013) proposal, which is well adapted to Gaussian priors such as (7). Noting that $\mathbb{F}(y_{1:K}\mid y_{0}=y)$ is a linear transformation of Gaussian variables $\eta$, and writing $\eta^{j}$ for those corresponding to $Y^{j}_{1:K}$, this consists in sampling $q$ as follows: given a step size $\delta>0$, (i) sample $\eta^{\prime}\sim\mathcal{N}(0,I)$, (ii) set $\eta^{*}=\frac{2}{2+\delta}\,\eta^{j}+\sqrt{1-\frac{4}{4+4\delta+\delta^{2}}}\,\eta^{\prime}$, and (iii) form $Y^{*}$ from $\eta^{*}$. Importantly, this can be done with zero memory cost if one has access to the random number generator that was used to sample $\eta^{j}$ originally, in which case we can simply re-simulate $\eta^{j}$ alongside $\eta^{\prime}$ at zero memory cost and negligible (compared to evaluating the SDE drift) computational cost.
Remark 3.3.
The parameter $\delta$ represents an exploration-exploitation trade-off: $\delta$ small corresponds to slow but likely successful exploration of the state-space, $\delta$ large corresponds to fast but likely unsuccessful exploration. In practice, for $\delta=0$ (which would result in a non-ergodic algorithm), the acceptance rate is fully dictated by the pseudo-marginal, in part by the variance of the normalising constant estimates, and in part by pseudo-marginal MCMC methods not recovering perfect samplers (Andrieu et al.,, 2018, Appendix F) For a large $\delta$, the acceptance will be fully dominated by the fact that the two normalising constants for two different paths $Y^{k}_{0:T}$ and $Y^{*}_{0:T}$ will be very different making the chain get stuck. The $\delta$ parameter interpolates between these two regimes, whereby we need to correlate $y_{0:T}$ and $y^{\prime}_{0:T}$ so that we can get close enough to the “ideal” scenario of the sampler being almost fully dominated by the pseudo-marginal behaviour.
@@ 4 Experiments
In this section, we conduct quantitative and qualitative experiments to validate our proposed method (i.e., Gibbs-CSMC in Sections 2.1 and 3.1 and PMCMC in Section 3.2) on synthetic and real data. We compare to related methods: the standard particle filter (PF, Trippe et al.,, 2023; Dou and Song,, 2024), twisted particle filter (TPF, Wu et al.,, 2023), and conditional score matching (CSGM, Song et al.,, 2021). In all experiments, particle filters use stratified resampling, and the conditional particle filter uses the conditional “killing” resampling of Karppinen et al., (2023) instead of the multinomial resampling in Algorithm 1. Unless otherwise mentioned, we repeat every experiment 100 times independently, reporting the mean, standard deviation, and/or quantile of their results. Our implementations are available at https://github.com/zgbkdlm/fbs.
4.1 SYNTHETIC CONDITIONAL SAMPLING
In our first ablation experiment, we evaluate the convergence of the conditioning methods to a tractable conditional distribution. Here, the target $\pi(x\mid y)$ is the posterior of a Gaussian process regression model with known ground truth. We use a linear forward noising process, so that the reverse SDE (2) is computed exactly, isolating the error coming from the conditioning methods only. For each method, we generate 10,000 posterior samples, and assess the errors using Kullback–Leibler (KL) divergence, Wasserstein–Bures (Bures) distance, and mean absolute errors of the marginal means and variances. Details are given in Appendix E.2.
Table 1: Errors statistics for the the conditional sampling problem of Section 4.1. The left and right panels show the errors when using 10 and 100 particles, respectively. The columns “Mean” and “Variance” show the errors in the marginal mean and variance of the posterior. The number after PMCMC is the PCN parameter $\delta$ applied. Since CSGM is computed exactly, it here serves as the reference and is not for comparisons.
@@ 4.1 SYNTHETIC CONDITIONAL SAMPLING
As shown by Table 4.1, Gibbs-CSMC outperforms all other methods for all metrics, effectively correcting the biase of the PF approach (Trippe et al.,, 2023). Our PMCMC approach also improves on PF but is sensitive to the calibration of the PCN parameter $\delta$: errors decrease when reducing $\delta=0.005$ to $\delta=0.001$. While PMCMC shows good marginal mean and variance errors, its KL performance suggests poor approximation to the off-diagonal GP posterior covariance. TPF performs worse than PF with higher standard deviations. Increasing the number of particles improves errors across all methods. More qualitative results are given in Appendix E.2.
Figure 1: Autocorrelations (of the worst dimension) in Section 4.1. For each of the 100 dimensions the autocorrelation is averaged over four chains. We see that Gibbs-CSMC outperforms PMCMC, with higher particle counts reducing autocorrelation.
Figure 1 shows the autocorrelations of our Gibbs-CSMC and PMCMC methods. Gibbs-CSMC exhibits better autocorrelation than PMCMC, even with fewer particles. PMCMC’s autocorrelation improves by decreasing the PCN parameter, mostly due to the variance of the log-likelihood estimation (see Section 3.2) of the PMCMC method (Andrieu and Roberts,, 2009; Doucet et al.,, 2015), sometimes preventing the algorithm from moving, and which is negatively affected by $\delta$. Lastly, the figure also confirms that increasing the number of particles improves the autocorrelations overall.
Figure 2: Conditional sampling errors on a Gaussian Schrödinger bridge: PF (ideal) uses the exact posterior (generally unavailable), while PF (approximate) uses a standard Normal. Gibbs sampling significantly improves quality, even over PF (ideal), underscoring PF’s inaccuracies.
@@ 4.2 CONDITIONAL SAMPLING ON NON-SEPARABLE NOISING PROCESSES
A key merit of our method is generalisability: it works with both separable and non-separable noising processes. Precisely, it does not need $X_{t}$ and $Y_{t}$ of the noising process in (4) to be independent given $X_{0}$ and $Y_{0}$. In contrast, standard PF approach of Dou and Song, (2024); Trippe et al., (2023) relies on separability. For non-separable processes, their method requires sampling $X_{0}$ to generate a path $t\mapsto Y_{t}$, ideally from the true conditional distribution, ideally from the true $X_{0}\sim\pi(x\mid y)$, but in practice, approximate samplers are used, introducing biases that are difficult to remove.
To quantify the bias from the non-separability, we perform conditional sampling on a Gaussian Schrödinger bridge model (for which closed-form expressions are available, see Bunne et al.,, 2023) using our Gibbs-CSMC and the PF of Trippe et al., (2023). We do not include PMCMC given its non-applicability here (see Section 3.2). Continuing with the GP regression problem from Section 4.1, we replace the reference $\mathcal{N}(0,I)$ with $\mathcal{N}(0,\Sigma)$, where $\Sigma$ is a random covariance matrix from a Wishart distribution, to highlight the impact of non-separability. Results in Figure 2 show that approximating the initial $\pi(x\mid y)$ leads to significant errors compared to the PF using exact initial sampling. Additionally, increasing the number of particles does not mitigate this initial bias, underscoring the shortcomings of the standard approach in non-separable models.
Figure 3: MNIST super-resolution (x4) on a non-separable Schrödinger bridge noising process. Each method shows two samples (more are in Figure 7). We find that PF is significantly affected by its initialisation of $X_{0}$, while Gibbs-CSMC is largely unaffected.
We also qualitatively assess the MNIST super-resolution on a pre-trained Schrödinger bridge. This time, we lack an exact $X_{0}$ sampler for the PF method, and we thus use a uniform random $X_{0}$, and an $X_{0}$ using linear interpolation of $y_{0}$. Results in Figure 3 show that the initial $X_{0}$ in the forward noising process significantly affects the outcomes. Using uniform $\mathcal{U}(0,1)$ initialisation leads to substantial deviations from the truth. In contrast, linear interpolation, a better initialisation, yields results that are closer to the truth, though the quality remains poor. Conversely, Gibbs-CSMC does not have the initialisation issue and recovers the original digit as the MCMC chain converges. More results are given in Appendix E.3.
@@ 4.3 INPAINTING AND SUPER-RESOLUTION
In this section, we apply our methods for inpainting and super-resolution on MNIST (resolution 28 by 28) and CelebA-HQ (resolution 64 by 64) images. For a complete description of the inpainting and super-resolution problems and additional experimental details, we refer the readers to Appendix E.4. Given pre-trained (via the standard score matching method of Song et al.,, 2021) unconditional generative diffusions for MNIST and Celeba-HQ images, we then perform conditional sampling of missing (inpainting) or deblurred (super-resolution) of images given an observation thereof. To quantify the results, for each test image we recover 100 conditional image samples, and we compute the peak signal-to-noise ratio (PSNR), structural similarity index (SSIM), and learnt perceptual patch similarity (LPIPS, Zhang et al.,, 2018) metrics averaged over the 100 samples.
@@
PF $15.85\pm 4.5$ $0.74\pm 0.2$ $10.14\pm 1.8$ $0.32\pm 0.2$ $17.34\pm 4.5$ $0.79\pm 0.1$ $11.66\pm 2.4$ $0.45\pm 0.2$
Gibbs-CSMC $16.79\pm 4.6$ $0.77\pm 0.1$ $10.84\pm 1.8$ $0.38\pm 0.2$ $17.82\pm 4.1$ $\mathbf{0.82}\pm 0.1$ $14.06\pm 3.4$ $0.61\pm 0.2$
PMCMC-0.005 $\mathbf{17.51}\pm 4.4$ $\mathbf{0.80}\pm 0.1$ $\mathbf{12.88}\pm 2.9$ $\mathbf{0.54}\pm 0.2$ $\mathbf{17.84}\pm 4.5$ $0.81\pm 0.1$ $\mathbf{14.22}\pm 2.8$ $\mathbf{0.63}\pm 0.2$
TPF $13.45\pm 3.1$ $0.48\pm 0.2$ $9.67\pm 1.6$ $0.19\pm 0.1$ $13.89\pm 3.2$ $0.51\pm 0.2$ $10.38\pm 2.0$ $0.25\pm 0.1$
CSGM $15.17\pm 4.2$ $0.71\pm 0.2$ $9.72\pm 1.6$ $0.28\pm 0.1$leftarrow $\leftarrow$ $\leftarrow$ $\leftarrow$
@@
PF $22.88\pm 3.7$ $0.85\pm 0.05$ $0.05\pm 0.03$ $23.29\pm 1.8$ $0.79\pm 0.05$ $0.11\pm 0.05$ $23.98\pm 3.6$ $0.870.05$ $\mathbf{0.04}\pm 0.02$ $24.57\pm 1.7$ $0.82\pm 0.04$ $0.09\pm 0.04$
Gibbs-CSMC $22.86\pm 3.7$ $0.86\pm 0.05$ $0.05\pm 0.03$ $23.75\pm 1.8$ $0.80\pm 0.05$ $0.11\pm 0.06$ $24.22\pm 3.5$ $0.87\pm 0.04$ $\mathbf{0.04}\pm 0.03$ $25.07\pm 1.8$ $0.84\pm 0.04$ $0.09\pm 0.05$
PMCMC-0.005 $\mathbf{23.96}\pm 3.5$ $\mathbf{0.87}\pm 0.05$ $\mathbf{0.04}\pm 0.03$ $\mathbf{24.37}\pm 1.7$ $\mathbf{0.82}\pm 0.04$ $\mathbf{0.10}\pm 0.04$ $\mathbf{24.68}\pm 3.5$ $\mathbf{0.88}\pm 0.04$ $\mathbf{0.04}\pm 0.02$ $\mathbf{25.30}\pm 1.8$ $\mathbf{0.85}\pm 0.04$ $\mathbf{0.07}\pm 0.03$
TPF $13.79\pm 2.4$ $0.59\pm 0.06$ $0.20\pm 0.07$ $12.71\pm 1.6$ $0.44\pm 0.08$ $0.23\pm 0.10$ $13.93\pm 2.4$ $0.58\pm 0.06$ $0.20\pm 0.07$ $12.92\pm 1.6$ $0.45\pm 0.08$ $0.23\pm 0.10$
CSGM $22.86\pm 3.6$ $0.85\pm 0.05$ $0.05\pm 0.03$ $23.48\pm 1.8$ $0.79\pm 0.05$ $0.11\pm 0.05$leftarrow $\leftarrow$ $\leftarrow$ $\leftarrow$ $\leftarrow$ $\leftarrow$
@@
In Figure 4, we illustrate the methods for inpainting and super-resolution on a single task. We see that PF, TPF, and CSGM can generate unrealistic samples, and the super-resolution samples are sometimes contaminated by pixel distortions (e.g., the isolated white dots in the second panel, first row, third-fifth columns in the figure). This is particularly true for TPF despite it largely higher computational cost (due to taking gradient with respect to the score function in its twisting function, Wu et al.,, 2023) as it requires to further approximate the likelihood model $\pi(y\mid x)$ by a smoother version to be usable. However, our Gibbs-CSMC and PMCMC methods are able to produce better quality samples, albeit more correlated ones, as those wrong samples are often rejected by the MCMC kernel. For more examples and results, we refer the readers to Appendix E.4.